OpenStudy (anonymous):

What's 6.75% as a fraction ?

OpenStudy (anonymous):

6.75/100.

OpenStudy (anonymous):

in simplest form my bad

OpenStudy (anonymous):

Simplest form: 27/400

OpenStudy (mathstudent55):

Percent literally means "per 100" 6.75 percent means 6.75/100

OpenStudy (stamp):

% = per cent = / 100 6.25% = 6.25 / 100 Multiply top and bottom by 4: 6.25(4) / 100(4) = 25 / 400 What do you get when you simplify this fraction?

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

@stamp, it's not 25/400, it's 27/400. She's asking for 6.75! not 6.25..

OpenStudy (stamp):

I notice that now. Good eyes catching my mistake :) \[27/400 = 3^3/2^45^2\] Does not reduce.